OpenStudy (anonymous):

Find the point on the line y = 9x closest to the point (1, 0).

OpenStudy (anonymous):

We can solve this two ways, first using calculus, then without: In both cases, first we simplify the equation of the line to 3x + y - 1 = 0 and then to y = 1 - 3x First way, using calculus: The distance from any point p on the line to (1,5) is given by sqrt ( (x-1)^2 + (y-5)^2 ) We wish to minimize that. We can ignore the sqrt, since for positive d, the minimum of d and minimum of d^2 occur for the same d. We substitute the value of y from the equation of the line into the formula of the distance, take the derivative, and find where it is 0. (x-1)^2 + (y-5)^2 (x-1)^2 + (1-3x-5)^2 x^2 - 2x + 1 is the square of the 1st term the second term = -3x-4 which has the same square as 3x+4, which is 9x^2 +24x +16 so we have the sum of the two: 10x^2 + 22x + 17 Derivative is 20x + 22 Setting that to 0, we get 20x = -22 and x = -11/10 y = 1 - 3x = 1 + 33/10 or So the point is (-1.1, 4.3) Second way: Another way to find it without using calculus, is to derive the line through 1,5, perpendicular to the first line. First line has slope -3, so new line will have slope 1/3 1/3 (1-x) = (5-y) 1/3 - x/3 = 5-y 1 - x = 15 - 3y 3y = x +14 Plugging in y from the first line, we get 3 (1 - 3x) = x + 14 3 - 9x = x +14 -10x = 11 x = -11/10 which agrees with the other solution.